if it takes 80.0ml of 1.500 M KOH to neutralize120.0ml of given H3PO4 solution, what is the molarity of this phosphoric acid solution?
1
Expert's answer
2018-11-07T14:55:09-0500

The molarity is 0.33 M.


The equation is:


3KOH + H3PO4 → K3PO4 + 3H2O


The number of moles of KOH = 1.5mol/L*0.08L = 0.12 mol


From the equation the number of moles of phosphoric acid must be 1/3 of this:


n(H3PO4) = 0.12/3 = 0.04 mol


Molarity = 0.04mol/0.12L = 0.33 M